OpenStudy (anonymous):

An empty truck with a mass of 2ton has a maximum acceleration of 1m/s . What will the maximum acceleration be when it carries a load of 1000kg

OpenStudy (anonymous):

With a maximum force being the product of a given mass and acceleration, and assuming the "ton" used here refers to a short ton (1 ton = 907kg): Mt = Mass of truck = 2 tons = 1814.37kg a = Acceleration = 1m/s^2 F = Maximum force = Mt * a = 1814.37kg * 1m/s^2 = 1814.37N Applying the maximum force to the truck with a 1000kg load, we have: Mn = Net mass = Mt + 1000kg = 2814.37kg F = Maximum force = 1814.37N a = F/Mn = 1814.37N/2814.37kg = .645m/s^2 Therefore, the maximum acceleration is 0.645m/s^2. Again, note that this calculation assumes you are using the short ton (1 ton = 907kg). Other versions include the metric ton (1 ton = 1000kg) and the long ton (1 ton = 1016kg).